Hydrogen peroxide in aquous solution decomposes on warming to give oxygen(under condition where one mole of gas occupies 24dm³).100 cm³ of`X` M solution of H2O2 produces 3dm³ of O2.thus X is??

Grade:12th pass

3 Answers

Ambarish Kumar Banerjee
7 years ago
from equation, 24 dm3 of o2 comes from 2 moles of h2o2.
                      3 dm3 of o2 comes from ¼ moles of h2o2
  100 cc=10^-1 dm3. therefore x/10=1/4. saupposing molarity is x)
          x=2.5M. please approve if its right
Rax
6 years ago
From equation 24 dm3 of o2 comes from 2 mol of h2o23dm3 of o2 comes from 1/4 moles of h2o2Now molarity is equal to(Mole/vol (ml))*1000Therefore x=(1/4*1000)/100 =2.5
Rax
6 years ago
From equation 24 dm3 of o2 comes from 2 mol of h2o2 3dm3 of o2 comes from 1/4 moles of h2o2 Now molarity = (Mole/vol (ml))*1000 Therefore x=(1/4*1000)/100. =2.5